OpenStudy (anonymous):

Divide as indicated. Be sure to reduce your answer to lowest terms. (The numerator and denominator of the answer should not have any factors in common.) (Problem in comments...)

OpenStudy (anonymous):

\[\frac{ 10x^2+35x+15 }{ x^2-16 }\div \frac{ 20x^2+40x+15 }{ 2x^2-5x-12 }\]

OpenStudy (anonymous):

@ganeshie8 \[\frac{ 5(2x^2+7x+3) }{ (x+4)(x-4) }\div \frac{ 5(4x^2+8x+3) }{ (2x+3)(x-4) }\]

OpenStudy (anonymous):

Now, I flip the last fraction

OpenStudy (anonymous):

No, I continue factoring, then I flip it.

OpenStudy (anonymous):

\[\frac{ x+3 }{ x+4 }\] That's my answer! :)

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)